# Probing the radiation-dominated regime of laser-plasma interaction in multi-beam configurations of petawatt lasers

### v1

@@ Abstract
We model numerically the ultrarelativistic dynamics of a dense plasma microtarget in a focus of several intersecting femtosecond laser pulses of multi-petawatt power each. The aim is to examine perspective future experimental approaches to the search of the Inverse Faraday Effect induced by radiation friction. We show that multi-beam configurations allow lowering the peak laser power required to generate a detectable quasi-static longitudinal magnetic field excited due to the radiation reaction force. The effect remains significant at angles around $10^{\rm o}$ between the beams, vanishes when the angle exceeds $20^{\rm o}$, and remains rather stable with respect to variations of relative phases and amplitudes of the beams. We conclude that using four infrared femtosecond linearly polarized pulses, 15 petawatt power each, crossing at angles $\approx 10^{\rm o}$, the radiation-dominated regime of laser-plasma interaction can be experimentally demonstrated.
@@ I Introduction
Effects of radiation friction or, equivalently, the radiation reaction force in dynamics of charged particles, beams and plasma subject to intense electromagnetic fields have remained in the scope of interest of strong-field physics for several latest decades. Recently physics of the radiation reaction force (RRF) has become even more topical, thanks to the considerable progress in design and construction of multi-petawatt laser sources potentially capable to produce ultrashort electromagnetic pulses of intensity $10^{23}$/cm2 and higher (see reviews [1, 2, 3] for information on the current status of high-power laser facilities). The theoretical framework for treatment of the RRF in classical electrodynamics has been amply developed [4, 5] in the past, and various signatures of radiation reaction in the dynamics and radiation of fast particles in external fields such as radiation damping in accelerators or radiation losses for cosmic rays had become tutorial examples. However, the subject remains topical and essentially intricate because of the two circumstances raising at the overlap of plasma physics and that of extreme electromagnetic fields. Firstly, in a dense plasma driven by extremely strong electromagnetic fields of short focused laser pulses, the RRF is expected to alter the collective dynamics providing conditions for a new interaction regime called radiation-dominated dynamics and enriched by new mechanisms of particle acceleration and magnetic fields generation. Secondly, applying intensities above $10^{24}-10^{25}$/cm2 or using pre-accelerated plasma interacting with pulses of lower intensity one may enter the quantum regime of radiation reaction, whose theoretical understanding remains incomplete [6]. Experiments in this intensity domain can provide fundamental knowledge useful for the development of a comprehensive quantum theory of radiation reaction. For the detailed description of this research field, which currently remains dominated by theory with just a few experimental demonstrations of RRF reported so far [7, 8, 6], we send the reader to reviews [9, 10, 1, 2, 3] and the literature quoted therein.
@@
The Inverse Faraday effect (IFE) [11, 12, 13] induced by the RRF in collisionless plasma interacting with circularly polarized (CP) electromagnetic radiation of extreme intensity can serve as an example of such collective plasma phenomenon arising due to radiation friction [14, 15, 16]. The IFE provides a new mechanism for the excitation of strong quasi-static magnetic fields in plasma, and, more importantly, it can be used as a benchmark of the radiation-dominated dynamics in classical or quantum regimes of laser-plasma interactions. It was shown in Ref.[14] that the synchrotron-like emission by fast electrons in a CP laser pulse results in a considerable irreversible angular momentum transfer from the field to the plasma followed by the excitation of ring-like electron currents in the polarization plane, which in turn generate a quasi-static magnetic field. At laser intensities $\simeq 10^{24}$/cm2, this magnetic field can reach an amplitude of several giga-Gauss (GG). More precisely, for the interaction of a CP pulse of $\lambda=800$nm wavelength and a super-Gaussian lateral intensity distribution of width $r_{0}=3.8\lambda$ the generation of a 1 GG magnetic field was predicted in the simulation [14] for the peak intensity $J=4.4\cdot 10^{23}$/cm2. The magnetic field exists for more than 100 fs in the volume $\simeq 50\lambda^{3}$ and should be accessible for detection by proton diagnostics.
@@
The aforementioned parameters require the peak laser intensity $\approx 80$PW at $\approx 1.2$kJ pulse energy. But despite the significant advances in petawatt (PW) laser technologies demonstrated in the two past decades, such parameters are still far from reach in laboratories. The peak power of 10-15 PW at a pulse duration 20-30 fs looks as an optimistic promise for the nearest future. The corresponding energies are 100-300 J per pulse. Several laser facilities [17, 18, 19, 20, 21] are filling currently the interval of 1-10 PW power. With this power level secured, projects of multibeam sub-exawatt laser installations will become realistic. The laser facility XCELS combining up to 12 synchronized 10-15 PW laser beams can serve an example of such projects [22], which can be put in realization within the nearest decade.
@@
The possibility to superimpose coherently several laser beams in the focal area will allow lowering the power and pulse energy thresholds required to achieve the radiation dominated regime, including the generation of strong magnetic fields through the IFE. The muti-beam scheme would also eliminate the necessity of converting an initially linearly polarized (LP) beam into a CP one as it will be sufficient to use several pairs of LP beams with mutually orthogonal polarizations propagating at small angles to each other. In this work, we examine such a scheme numerically by modeling the laser-plasma interaction within the particle-in-cell (PIC) method using the codes UMKA [23, 24, 25] and SMILEI [26]. We aim at the following objectives: (a) to estimate the minimal laser power and pulse energy required for the generation of a significant, detectable magnetic field due to the radiation reaction effect in a multi-beam configuration and (b) to check the sensibility of the generated magnetic field peak amplitude and spatial distribution to the relative phase shifts of the beams, variations in their amplitudes and to the angle at which the beams intersect. The robustness of the scheme with respect to the first two factors is of particular importance, since they are difficult to control in multi-PW sources with a low repetition rate. The angle between the beams’ propagation directions needs optimization because the electromagnetic field structure close to CP vanishes quickly with this angle increasing, while using nearly parallel beams intersecting at small angles is technically difficult. The output of our study will help to formulate requirements to the architecture of multi-beam multi-PW laser source XCELS and other laser facilities of similar class.
@@
The structure of the paper is following. In the next section II we formulate the statement of the problem and discuss parameters of the laser pulses and plasma. Section III is dedicated to presentation and discussion of numerical results. Section IV presents conclusions and outlook. Preliminary results of our calculations related to the two-beam scheme were reported in [27].
@@ II.1 Interaction scheme and parameters
with $c$ being the speed of light. A position-dependent analog $a(\bf r)$ of (2) is defined by replacing the field amplitude $E_{0}$ with $E(\mathbf{r})=\sqrt{<E^{2}>(\mathbf{r})}$, where <>expectation<...>< … > denotes time averaging. The plasma parameters are chosen to keep them close to a set of earlier calculations [14, 15, 16], where we studied the IFE in a single CP beam of intensity $J\approx 10^{24}$/cm2 and higher. For LP pulses of frequency $\omega=2.4\cdot 10^{15}$s-1, $a_{0}=1$ at $J=2.1\cdot 10^{18}$/cm2. In a single CP laser beam, our previous calculations [14, 15] did not reveal any significant longitudinal magnetic field for $a_{0}<200$. Thus in the present calculation we are interested in intensities $\approx 8\cdot 10^{22}$/cm2 in a single beam.
We consider laser pulses ejected through the left boundary of the simulation box. This is achieved through a time-dependent boundary condition set on the $x=-5\lambda$ plane. The finite-difference time-domain method [28, 29] with the total-field/scattered field formulation [30] in x-direction is used to propagate the electromagnetic fields. To obtain the distributions of the magnetic field strength $\mathbf{B}_{\perp}$ on the left boundary we consider a super-Gaussian beam  [^1] Note that the functions in (3) do not solve Maxwell equations, even in the paraxial approximation. Thus, this formula is only used to determine the time dependent boundary condition, while the electromagnetic field in the interaction volume is found by solving the Maxwell equations numerically. of the following form
$\begin{split}\mathbf{B}_{\perp}=E_{0}\mathbf{p}\exp&\left\{-\frac{(y^{\prime 2}+z^{\prime 2})^{2}}{r_{0}^{4}}-\frac{(x^{\prime}-ct)^{4}}{r_{c}^{4}}\right\}\times\\ &\times\left(\begin{array}[]{ll}\cos(kx^{\prime}-\omega t)\\ \sin(kx^{\prime}-\omega t)\end{array}\right).\end{split}$ (3)
@@
Here $E_{0}$ is the electric and magnetic peak field amplitude in (2) and $\mathbf{p}$ is the field polarization vector. Coordinates $(x^{\prime},y^{\prime},z^{\prime})$ in (3) are introduced for each beam individually such that the wave vectors ${\bf k}^{\prime}=(k,0,0)$ are orthogonal to the corresponding $(y^{\prime},z^{\prime})$ planes. Primed coordinates are related to laboratory ones used in the simulation by rotation through an angle $\theta$ in the $(x,y)$ plane for the two-beam setup and in the $(x,y)$ and $(x,z)$ planes for the four-beam setup. Schematic representation of both setups is shown on Fig. 1(a,b). The beam lateral waists and the longitudinal pulse lengths are taken $r_{0}=3.8\lambda$ and $r_{c}=3\lambda$ correspondingly. The effective pulse duration is $\tau=3\lambda/c=6\pi/\omega\approx 7.2$ fs. This is roughly 3 times shorter than a typical pulse duration for most of the multi-PW sources. This difference is not of crucial importance for our analysis, while the application of such short pulses considerably accelerates demanding numerical simulations.
@@
The applied boundary condition (3) does not focus the field on the plasma surface, instead the beams appear at $x=0$ defocused compared to the plane $x=-5\lambda$. This defocusing is relatively weak thanks to small distance between the boundary and the plasma surface. Thus the peak intensity estimated using of the field distribution (3) on the boundary does not differ much from the value reached on the plasma slab surface. Solutions found in free space (without plasma) allow to check, how quickly the field in the area where the beams overlap starts deviating from CP with the angle $\theta$ increasing. We characterize this deviation by the distributions in dimensionless intensity ${\tilde{J}}=a_{0}^{2}({\bf r})$ and in the density of x-projection of angular momentum ${\bf m}({\bf r})=[{\bf r}[{\bf E}{\bf B}]]/4\pi c$ shown in Fig. 1(d-g). For a CP beam the $a^{2}$-distribution is smooth (Fig. 1(d)), and the $m_{x}$ density is everywhere positive, while for $\theta\approx 10^{\rm o}$ a considerable difference is already clearly seen. The sign of the integrated projection $M_{x}$ flips at $2\theta\approx 10^{\rm o}$ (Fig. 1(c)). The time dependence of the electric field projection $E_{x}$ and of the absolute value of its lateral component $E_{\bot}=\sqrt{E_{y}^{2}+E_{z}^{2}}$ shown in panels of Fig. 1(h-k) also demonstrates the deviation from CP increasing with the growth of the convergence angle. For $2\theta>10^{\rm o}$ the field structure remains close to CP on the axis $y=z=0$, while at $y=4\lambda$ the deviation is apparent. Taking into account that the IFE is induced by electron currents excited near the laser beam lateral edge ($\approx 3-4\lambda$ in our case) [14] this deviation should considerably suppress the magnetic field generation. Our further simulations (not present in the plot) show that the field structure in the overlap area of the beams depends on the boundary condition. In particular, a Gaussian boundary condition such that the beams overlap in their focal waists leads to a slower decrease in the degree of CP with $\theta$ increasing. However, this does not change our results qualitatively. Summarizing the information delivered in Fig. 1, we conclude that with $\theta$ increasing, the field structure in the focus deviates form CP rather quickly so that the case $\theta>10^{\rm o}$ is not of interest for the detection of IFE signatures [27].
@@
Numerical runs described in the next subsection were arranged such that the time instant $t=0$ is that one when the laser pulse front edge arrives to the front surface of the plasma slab at $x=0$ so that $t>0$ is time past after the start of the interaction. With the efficient pulse duration $\tau=3T_{L}\approx 7.2$fs the laser field is being entirely reflected by the plasma during $\tau_{\rm int}\approx 15-20$fs, which can be considered as the full interaction time. We look at the magnetic field excited in the plasma after the interaction, typically at $t=30-40T_{L}$. For the super-Gaussian pulse (3) the peak power is
@@
with $J_{0}=2.1\cdot 10^{18}{\rm W/cm}^{2}$ for $\lambda=800$nm and LP. Taking $a_{0}=200$ one obtains from (4) $P\approx 16$PW, the value close to that expected at the XCELS laser facility under development [22]. The full pulse energy is about 150W. Note that (4) only gives an estimate of the peak laser power, because the field structure in the interaction volume obtained through the numerical solution of the Maxwell equations deviates considerably from (3).

### v2

@@ Abstract
We model numerically the ultrarelativistic dynamics of a dense plasma microtarget in a focus of several intersecting femtosecond laser pulses of multi-petawatt power each. The aim is to examine perspective future experimental approaches to the search of the Inverse Faraday Effect induced by radiation friction. We show that multi-beam configurations allow lowering the single beam peak laser power required to generate a detectable quasi-static longitudinal magnetic field excited due to the radiation reaction force. The effect is significant at angles around $10^{\rm o}$ between the beam propagation axes, almost vanishes when the angle exceeds $20^{\rm o}$, and remains rather stable with respect to variations of relative phases and amplitudes of the beams. Quantum recoil accounted within semi-classical approach is shown to considerably suppress the longitudinal magnetic field, which however remains sizable. We conclude that using four infrared femtosecond linearly polarized pulses, 15 petawatt power each, crossing at angles $\approx 10^{\rm o}$, the radiation-dominated regime of laser-plasma interaction can be experimentally demonstrated.
@@ I Introduction
Effects of radiation friction or, equivalently, the radiation reaction force in dynamics of charged particles, beams and plasma subject to intense electromagnetic fields have remained in the scope of interest of strong-field physics for several latest decades. Recently physics of the radiation reaction force (RRF) has become even more topical, thanks to the considerable progress in design and construction of multi-petawatt laser sources potentially capable to produce ultrashort electromagnetic pulses of intensity $10^{23}$/cm2 and higher (see reviews [1, 2, 3] for information on the current status of high-power laser facilities). The theoretical framework for treatment of the RRF in classical electrodynamics has been amply developed [4, 5] in the past, and various signatures of radiation reaction in the dynamics and radiation of relativistic particles in external fields such as radiation damping in accelerators or radiation losses for cosmic rays had become tutorial examples. Recent experimental advances in the observation of interactions of ultrafast charged particles with crystals in the regime of channeling have opened the way to detect effects of quantum radiation reaction on the motion of positrons due to strong electric fields in aligned crystals [6, 7]; see also [8, 9] for the theoretical treatment. This research has set an important milestone showing clear signatures of the quantum radiation reaction regime. Together with the recent experiments on the interaction of laser-accelerated electron bunches with strong counter-propagating laser pulses [10, 11], these studies have considerably extended knowledge of the realm of the quantum interaction regime, although its theoretical treatments still remain incomplete, and the transition from the fully quantum to the classical description is waiting for the ultimate clarification [12, 6].
Despite the aforementioned recent experimental and theoretical advances, the subject remains topical and essentially intricate because of the two circumstances raising at the overlap of plasma physics and that of extreme electromagnetic fields. Firstly, in a dense plasma driven by extremely strong electromagnetic fields of short focused laser pulses, the RRF is expected to significantly alter the collective dynamics providing conditions for a new interaction regime called radiation-dominated plasma dynamics and enriched by new mechanisms of particle acceleration and quasi-static magnetic fields generation. Secondly, applying intensities above $10^{24}-10^{25}$/cm2 or using pre-accelerated plasma interacting with pulses of lower intensity one may enter the multi-particle quantum regime of radiation reaction characterized by the development of quantum cascades and the creation of dense electron-positron plasmas, see e.g. [13, 14]. Experiments in this intensity domain can provide fundamental knowledge useful for the development of a comprehensive quantum theory of radiation reaction. For the detailed description of this research field we send the reader to reviews [15, 16, 1, 2, 3] and the literature quoted therein.
@@
The Inverse Faraday Effect (IFE) [17, 18, 19] induced by the RRF in collisionless plasma interacting with circularly polarized (CP) electromagnetic radiation of extreme intensity can serve as an example of the aforementioned collective plasma phenomena arising due to radiation friction [20, 21, 22]. The IFE provides a new mechanism for the excitation of strong quasi-static magnetic fields in plasma, and, more importantly, it can be used as a benchmark of the radiation-dominated dynamics in classical or quantum regimes of laser-plasma interactions. It was shown in Ref.[20] that the synchrotron-like emission by fast electrons in a CP laser pulse results in a considerable irreversible angular momentum transfer from the laser field to the plasma followed by the excitation of ring-like electron currents in the polarization plane, which in turn generate a quasi-static magnetic field. At laser intensities $\simeq 10^{24}$/cm2, this magnetic field can reach an amplitude of several gigagauss (GG). More precisely, for the interaction of a CP pulse of $\lambda=800$nm wavelength and a super-Gaussian lateral intensity distribution of width $r_{0}=3.8\lambda$ the generation of a 1 GG magnetic field was predicted in the simulation [20] for the peak intensity $J=4.4\cdot 10^{23}$/cm2. The magnetic field exists for more than 100 fs in the volume $\simeq 50\lambda^{3}$ and should be accessible for detection by proton diagnostics. Further calculations made with a semi-classical account of the quantum recoil effect on the dynamics and radiation of plasma electrons [22], showed, for these interaction parameters, considerable suppression of the IFE without, however, qualitative changes in the picture.
@@
The aforementioned interaction parameters can be reached at the peak laser power $\approx 80$PW and $\approx 1.2$kJ pulse energy. However, despite the significant advances in petawatt (PW) laser technologies demonstrated in the two past decades, such values are still far from reach in laboratories. The peak power of 10-15 PW at a pulse duration 20-30 fs looks as an optimistic promise for the nearest future. The corresponding energies are 100-300 J per pulse. Several laser facilities [23, 24, 25, 26, 27] are filling currently the interval of 1-10 PW power. With this power level secured, projects of multibeam sub-exawatt laser installations will become realistic. The laser facility XCELS combining up to 12 synchronized 10-15 PW laser beams can serve an example of such projects [28], which can be put in realization within the nearest decade. Note that presently several possible multibeam configurations remain under discussion for the design of the future XCELS installation. Apart of the mentioned $12\times 15$PW scheme, there is another, which involves fewer beams 50PW each.
@@
The possibility to superimpose coherently several laser beams in the focal area will allow lowering the single pulse energy and power thresholds required to achieve the radiation dominated regime, including the generation of strong magnetic fields through the IFE. The muti-beam scheme would also eliminate the necessity of converting an initially linearly polarized (LP) beam into a CP one as it will be sufficient to use several pairs of LP beams with mutually orthogonal polarizations and an appropriate phase shifts propagating at small angles to each other. In this work, we examine such a scheme numerically by modeling the laser-plasma interaction within the particle-in-cell (PIC) method using the codes UMKA [29, 30, 31] and SMILEI [32]. The former allows for simulating the plasma dynamics both in the classical and quantum regimes of interaction with the effect of photon recoil accounted for semi-classically in the second case along the method of Refs.[33, 34]. We aim at the following objectives: (a) to estimate the minimal single pulse energy and power required for the generation of a significant, detectable magnetic field due to the radiation reaction effect in a multi-beam configuration and (b) to check the sensibility of the generated magnetic field peak amplitude and spatial distribution to the relative phase shifts of the beams, variations in their amplitudes and to the angle at which the beams intersect. The robustness of the scheme with respect to the first two factors is of particular importance, since they are difficult to control in multi-PW sources with a low repetition rate. The angle between the beams’ propagation directions needs optimization because the electromagnetic field structure close to CP vanishes quickly with this angle increasing, while using nearly parallel beams intersecting at small angles is technically difficult. The output of our study will help to formulate requirements to the architecture of multi-beam multi-PW laser source XCELS and other laser facilities of similar class.
@@
The structure of the paper is following. In the next section we formulate the statement of the problem and discuss parameters of the laser pulses and plasma. Section III is dedicated to presentation and discussion of numerical results. Section IV reports conclusions and outlook. Preliminary results of our calculations related to the two-beam scheme were reported in [35].
@@ II.1 Interaction scheme and parameters
with $c$ being the speed of light. A position-dependent analogue $a(\bf r)$ of (2) is defined by replacing the field amplitude $E_{0}$ with $E(\mathbf{r})=\sqrt{<E^{2}>(\mathbf{r})}$, where <>expectation<...>< … > denotes time averaging. The plasma parameters are chosen to keep them close to a set of earlier calculations [20, 21, 22], where the IFE in a single CP beam of intensity $J\approx 10^{24}$/cm2 and higher was studied. For LP pulses of frequency $\omega=2.4\cdot 10^{15}$s-1, $a_{0}=1$ at $J=2.1\cdot 10^{18}$/cm2. In a single CP laser beam, our previous calculations [20, 21] did not reveal any significant longitudinal magnetic field for $a_{0}<200$. Thus in the present calculation we investigate the case of intensities $\approx 8\cdot 10^{22}$/cm2 in a single beam.
The laser pulse is initialized by setting a non-stationary boundary condition on the left boundary $x=x_{0}=-5\lambda$ of the computation domain. The electromagnetic field inside this domain is found by solving the Maxwell’s equations numerically. To set the boundary condition, we use super-Gaussian beams of the following shape
$f(x^{\prime},y^{\prime},z^{\prime},t)=\exp\left\{-\frac{(y^{\prime 2}+z^{\prime 2})^{2}}{r_{0}^{4}}-\frac{(x^{\prime}-ct)^{4}}{r_{c}^{4}}\right\}$ (3)
@@
The finite-difference time-domain method [36, 37] with the total-field/scattered field formulation [38] in x-direction is used to propagate the electromagnetic fields. Coordinates $(x^{\prime},y^{\prime},z^{\prime})$ in (3) are introduced for each beam individually such that the wave vectors ${\bf k}^{\prime}=~{}(k,0,0)$ are orthogonal to the corresponding $(y^{\prime},z^{\prime})$ planes. Primed coordinates are related to laboratory ones used in the simulation by rotation in an angle $\theta$ in the $(x,y)$ plane for the two-beam setup and in the $(x,y)$ and $(x,z)$ planes for the four-beam setup. Schematic representation of both setups is shown on Fig. 1(a,b). The beam lateral waists and the longitudinal pulse lengths are taken $r_{0}=3.8\lambda$ and $r_{c}=3\lambda$ correspondingly. The effective pulse duration is $\tau=3\lambda/c=6\pi/\omega\approx 7.2$ fs. This is roughly three times shorter than a typical pulse duration for most of the multi-PW sources. This difference is not of crucial importance for our analysis, while the application of such short pulses considerably accelerates demanding numerical simulations.
@@
The boundary condition (3) does not focus the field on the plasma surface, instead the beams appear at $x=0$ defocused compared to the plane $x=-5\lambda$. This defocusing is relatively weak thanks to a small distance between the boundary and the plasma surface. Thus the peak intensity estimated using of the field distribution (3) on the boundary does not differ much from the value reached on the plasma slab surface. Solutions found in free space (without plasma) allow to check, how quickly the field in the area where the beams overlap deviate from CP with the angle $\theta$ increasing. We characterize this deviation by the distributions in dimensionless intensity ${\tilde{J}}=a_{0}^{2}({\bf r})$ and in the density of x-projection of angular momentum ${\bf m}({\bf r})=[{\bf r}[{\bf E}{\bf B}]]/4\pi c$ shown in Fig. 1(d-g). For a CP beam the $a^{2}$-distribution is smooth (Fig. 1(d)), and the $m_{x}$ density is everywhere positive, while for $\theta\approx 10^{\rm o}$ a considerable difference from the CP case is already clearly seen. The sign of the integrated projection of the angular momentum of the field $M_{x}$ flips at $2\theta\approx 8^{\rm o}$ (Fig. 1(c)). The time dependence of the electric field projection $E_{x}$ and of the absolute value of its lateral component $E_{\bot}=\sqrt{E_{y}^{2}+E_{z}^{2}}$ shown in panels of Fig. 1(h-k) also demonstrates the deviation from CP increasing with the growth of the convergence angle. For $2\theta>10^{\rm o}$ the field structure remains close to CP on the axis $y=z=0$, while at $y=4\lambda$ the deviation is apparent. Taking into account that the IFE is induced by electron currents excited near the laser beam lateral edge ($\approx 3-4\lambda$ in our case) [20] this deviation should considerably suppress the magnetic field generation. Our further simulations (not present in the plot) show that the field structure in the overlap area of the beams depends on the boundary condition. In particular, a Gaussian boundary condition defined such that the beams overlap in their focal waists leads to a slower decrease in the degree of CP with $\theta$ increasing. However, this does not change our results qualitatively. Summarizing the information delivered in Fig. 1, we conclude that with $\theta$ increasing, the field structure in the focus deviates form CP rather quickly so that the case $\theta>10^{\rm o}$ is not of interest for the detection of IFE signatures [35].
@@
Numerical runs described in the next subsection were arranged such that the laser pulse front edge arrives to the front surface of the plasma slab $x=0$ at $t=0$ so that $t>0$ is time past after the start of the interaction. With the efficient pulse duration $\tau=3T_{L}\approx 7.2$fs the laser field is being entirely reflected by the plasma during $\tau_{\rm int}\approx 15-20$fs, which can be considered as the full interaction time. We look at the magnetic field excited in the plasma after the interaction, typically at $t=30-40T_{L}$. For the super-Gaussian pulse (3) the peak power is
@@
with $J_{0}=2.1\cdot 10^{18}{\rm W/cm}^{2}$ for $\lambda=800$nm and LP. Taking $a_{0}=200$ one obtains from (4) $P\approx 16$PW, the value close to that expected at the XCELS laser facility under development [28]. The full pulse energy is about 150J. Note that (4) only gives an estimate of the peak laser power, because the field structure in the interaction volume obtained through the numerical solution of the Maxwell equations deviates from the analytic form (3).